OpenStudy (anonymous):

Evaluate the function. f(x)=5^(ln(x+1))

OpenStudy (anonymous):

Is this the complete question? Is any other information given?

OpenStudy (anonymous):

oh whoops. Find f'(0). I know the answer is f'(0)=ln5. But I don't understand the process.

OpenStudy (anonymous):

\[f \left( x \right)=5^{\ln \left( x+1 \right)},\ln f \left( x \right)=\ln 5^{\ln \left( x+1 \right)}=\ln \left( x+1 \right)\ln 5\] diff. w.r.t.x \[\frac{ f'\left( x \right) }{ f \left( x \right) }=\frac{ \ln 5 }{ x+1 }\] \[f'\left( x \right)=f \left( x \right)\frac{ \ln 5 }{ x+1 }=5^{\ln \left( x+1 \right)}\frac{ \ln 5 }{ x+1 }\] i think now you can solve.

OpenStudy (anonymous):

Yes, thank you!

OpenStudy (anonymous):

you are welcome.

OpenStudy (anonymous):

Heyyy Im confused. What does diff. w.r.t.x mean?

OpenStudy (anonymous):

with respect to x